OpenStudy (babtaooche):

Help, I'll fan and medal How can x + 8 = −x + 7 be set up as a system of equations?

OpenStudy (babtaooche):

Answers: y = x + 8 y = −x + 7 y = x + 8 y = −3x + 7 y = x + 5 y = −2x + 11 y = 8 y = −3x + 7

OpenStudy (retireed):

Break up the first equation into two equations with both halves equal to y.

OpenStudy (retireed):

Like this...